Your firm currently has 1,187,810 shares of stock outstanding that sell for 36.22. Additionally, you've issued 32,812 bonds, with a par value of S1,000. They currently sell for 85.5% of par.

What is the firm's equity weight?

(Enter your answer as a percentage to 1 decimal place. ex: 12.3)
Answer #1

Total value of:

Equity=1,187,810*36.22=$43022478.2

Bonds=32,812*(1000*85.5%)=$28054260

Total value=(43022478.2+28054260)=$71076738.2

Hence equity weight=(43022478.2/71076738.2)

=60.5%(Approx).